OpenStudy (anonymous):

Find the value of (a) for which (x-2) is a factor of 3x^3 + ax^2 + x - 2

OpenStudy (anonymous):

\[3x^3 + ax^2 + x - 2=(x-2)(Ax^2+Bx+C)\]

OpenStudy (anonymous):

3x^3+ax^2+x-2=0 Since x-2 is a factor we can apply x=2 in the equation above. therefore 3(2^3)+a(2^2)+2-2=0 3(8)+4a=0 4a=-24 a=-6

OpenStudy (anonymous):

Ah, Ok. Thanks :)

OpenStudy (anonymous):

You are welcome.